A commercial jet travels 1560 mi in the same amount of time it takes a corporate jet to travel 1200 mi. The rate of the commercial jet is 120 mi/h greater than the rate of the corporate jet. Find the rate of the corporate jet.
Answer by lwsshak3(11628)

A commercial jet travels 1560 mi in the same amount of time it takes a corporate jet to travel 1200 mi. The rate of the commercial jet is 120 mi/h greater than the rate of the corporate jet. Find the rate of the corporate jet
let x=rate of corporate jet
x+120=rate of commercial jet
Travel time=distance/rate
1560/(x+120)=1200/x
1560x=1200x+144000
360x=144000
x=400 mi/hr
rate of corporate jet=400 mi/hr